Pablo invested $445,000 in two bonds. The first bond paid 3% interest per year and the second 8% per year. The total interest pablo received in one year was $2850. How much did he invest in each bond?
Answer by checkley77(12569)

.08x+.03(45,000-x)=2,850
.08x+1,350-.03x=2,850
.05x=2850-1,350
.05x=1,500
x=1,500/.05
x=30,000 A VALID ANSWER.
PROOF:
.08*30,000+.03(45,000-30,000)=2,850
2,400+.03*15,000=2,850
2,400+450=2,850
2,850=2,850